OpenStudy (anonymous):

If two adults and three children go to the movies, it costs $31.50. If one adult and five children go to the movies, it costs $35.00. What is the cost of one adult movie ticket?

OpenStudy (anonymous):

\[2a+3c=31.50\]\[a+5c=35.00\] Lets solve for a in the second equation since it seems to be the least complicated.\[a+5c=35\]\[a=35-5c\] Now that we know the value of a in terms of c we can solve for c. \[2(35-5c)+3c=31.50\]\[70-10c+3c=31.50\]\[70-7c=31.50\]\[7c=-38.5\]\[c=-5.50\] Now that we know the price of a children's ticket, we can calculate the amount of an adult's ticket by plugging 5.50 into one of the equations and solving for a. \[a+5(5.50)=35\]\[a+27.50=35\]\[a=35-27.5\]\[a=7.50\]Now we'll double check that answer by plugging it into the first equation. \[2(7.50)+3(5.50)=31.50\] \[15+16.5=31.50\]\[31.50=31.50\] Check. So an adult ticket costs $7.50.